Problem 1. Suppose we have the following data about a simple economy:

C = 10 + 0.75Yd

I = 50

G = T = 20

 (a) Find out the equilibrium level of national income.

(b) What is the size of the multiplier?




1
Expert's answer
2022-01-17T12:40:46-0500

 (a) The equilibrium level of national income.

At equilibrium; AD = AS. Therefore: Y-t = C+I+G

Y - 20 = 10+0.75Yd+50+20

Y - 20 = 0.75Yd + 80

Y - 0.75Yd = 100

0.25Y* = 100

Y = 400

(b) What is the size of the multiplier?

"multiplier = \frac{1}{1-MPC}"


MPC= 0.75

="\frac{1}{1-0.75} =4"


Multiplier = 4